# ExAMPC: the Data-Driven Explainable and Approximate NMPC with Physical Insights

### v1

@@ III-B MultiOutput Regression Trees, Neural Networks and Symbolic Regression
[…]
III-C One the use of AI for approximate NMPC
We focus on learning two types of output as in Figure 1:
  1. 1.
    the optimized NMPC OL trajectory solution at every instance $t$, for given state estimate $x_{0}(t)$: $\{x_{0}(t),\eta,p(t)\}\rightarrow\{\mathbf{x}_{(\cdot)}(t),\mathbf{u}_{(\cdot)}(t)\}$. The autonomous CL system evolves under uncertainties $\xi$ and scenario parameters $p(t)$ (such as target states, boundary conditions, etc..).
  2. 2.
Existing approaches like [7] approximate the NMPC solution to predict discrete sequences $\mathbf{x}_{(\cdot)}(t)=\{x_{0}(t),x_{1}(t+T_{s}),\dots,x_{N}(t+NT_{s})\}$ over the NMPC horizon length $N$ with a step size $T_{s}$, but face several limitations:
  • Data inefficiency: trajectories resulting from discrete methods such as direct multiple shooting scale poorly with horizon length and step size. While technologies involving transformers [7, 9] can handle longer trajectories, their data requirements limit scalability under changing CL conditions.
  • High output dimensionality for discrete sequences.
  • Lack of CTCP, potentially compromising safety.
  • Limited output explainability: individual sequence elements provide minimal insights into trajectory trends, dynamics, and sensitivity to input features and parameters, making them non-intuitive (e.g. factors affecting high velocity rates and accelerations).
The authors of [5] address the first three limitation by embedding the infinite dimensional continuous sequence into a finite set of B-spline coefficients, which have a convex hull property. However, B-splines are complex to construct, highly sensitive to the chosen knot sequence, and the coefficients offer limited physical insights on the trajectory trend and dynamics. Moreover, modeling time-series with high accuracy using B-splines requires a rather dense knot sequence, where poorly chosen knots can lead to ill-conditioning in the fitting problem.
RESAFE/COL [16] overcomes these B-spline limitations by using spectral collocation with orthogonal basis polynomials to achieve high solution accuracy and allows to impose continuous-time nonlinear constraint satisfaction by relying on a linear mapping between the coefficients that embed a Legendre-Spline and its extrema as in (4). This approach provides interpretable coefficients sequence revealing physical evolution information on the trajectory through the zero-order bias term $\alpha_{0}$ or the $i^{th}$ derivative terms. Moreover, it benefits from a better numerical conditioning and a low-order fitting: as the order $M$ increases, the high order-terms naturally vanish to zero if they do not increase the accuracy as the coefficients $\alpha_{j}$ decay faster than any polynomial in $j$ [18]. We refer to this method as a low-order embedding.
III-D Data generation: autonomous driving and racing
We demonstrate our work on autonomous driving and racing control applications. The NMPC is implemented in CasADi [19] using an SQP method with OSQP as the underlying QP solver, and the OCP is transcribed into an Nonlinear Programming Problem (NLP) using RESAFE/COL [16]. The NMPC uses a fused kinematic-dynamics bicycle model with a Pacejka tire formulation. The verification uses 15 Degrees-of-Freedom high-fidelity DTs of the vehicles in Simcenter Amesim. Two demonstration setups are created:
@@ III-D Data generation: autonomous driving and racing
Both setups integrate the NMPC as a standalone C-code library for co-simulation with the DTs with Simcenter Amesim for vehicle dynamics and Simcenter Prescan for environment simulation and sensor modeling for obstacle detection of crossing pedestrians and road users as visualized in Figure 7. The NMPC serves as the lowest-level control, executing trajectories from a high-level planner through steering, brake and throttle commands. The collected data is divided into training (64%), validation (16%) and testing (20%). Finally, the data is normalized per feature and per output to $[-1,1]$.
@@ IV AI as an explainable NMPC approximation
IV-A RESAFE/COL: a physics-informed NMPC approximation
The NMPC solves for continuous-time trajectories that are embedded in form of the coefficients of Legendre-Splines (3). As explained in Sec. III-C, the use of a Legndre-Spline with orthogonal basis offers two key advantages: 1) naturally regularized low-order embedding through decoupled and indepdent coefficients from each other, and 2) physical constraint enforcement via a linear coefficient mapping, enabling CTCP without discrete sampling of the time-series trajectory.
As illustrated in Figure 1, we collect CL data of the NMPC controlling a high-fidelity Digital Twin (DT) to train ML regression models to imitate the NMPC’s OL solution. Using similar input parameters being fed into the NMPC, the trained regression model would Learn2Optimize. Unlike standard Imitation Learning approaches that learn only the NMPC’s policy or first control action $u(t=t_{0})$, we propose to learn the complete time-series of this policy evolution $u(t)$, and that of the states $x(t)$. This approach would capture the optimization framework linking between predicted states and control actions. It also enables an effective warm-starting strategy for the NMPC which is known to have benefits on the numerical efficiency and helps speeding-up the computation in methods such as SQP. A baseline approach to this OL trajectory regression predicts the sequence of the sampled discrete trajectory points, or the embedding coefficients using a Mean Squared Error (MSE) loss:
$\mathcal{L}=\mathcal{L}_{MSE}=\sum_{i=1}^{N_{batch}}\lVert(\bar{\bm{\alpha}}_{i}-\bm{\alpha}_{i})\rVert^{2}/N_{batch}$ (5)
@@ IV-A RESAFE/COL: a physics-informed NMPC approximation
where $\epsilon_{tol}$ defines the violation threshold tolerance. Note that linear state and control constraints of the form $\underline{x}\leq x(t)\leq\overline{x}$ are expressed in the generic form of $g(x,u)\leq 0$ for conciseness. An example of the extrema elements of the convex hulls $\mathcal{P}_{u}$ are shown in shaded yellow area in Figure 1 over the decoded control action trajectories.
@@ IV-B Multistep prediction using coefficients: a data efficient and explainable approach
The proposed Legendre-Spline embedding addresses the challenges stated in Sec. III-C by encoding physical information through the coefficients: the zero order coefficient $\alpha_{0}$ for the trajectory bias or mean term, the first order coefficient $\alpha_{1}$ for information about the rate of change, the second-order coefficient $\alpha_{2}$ about acceleration characteristics and the higher-order terms on additional dynamic features of the trajectory. This representation enables engineers to interpret and shape the OCP’s OL behavior through physically meaningful parameters that can be explained using SHAP. Therefore, by combining (3) with (6) and XAI tools, we allow multistep sequence prediction in one shot while requiring little data due to the low dimensionality of the prediction, and while having physical insights as shown in the framework of Figure 1.
@@ IV-C ExAMPC as a warm-starter for NMPC
We train two RNNs to approximate the NMPC for the AVP use case: one using the baseline with MSE (MSE-RNN) and another with the RESAFE/COL-type of loss as in (6) for CTCP with $\gamma=1$ (RESAFE-RNN). As shown in Figure 2, the approximate NMPC as RESAFE-RNN achieves 556 continuous-time constraint violations out of 57632 testing instances. The total loss is equal to 2.05e-04 divided into a MSE of 2.0e-04 and CTCP of 4.9e-06. In contrast, the baseline MSE-RNN using only coefficients learning results in 8113 violations out of 57632 instances, with a MSE on the coefficients of 1.7e-04 but a CTCP of 8.5e-03. Overall, the RESAFE/COL approach with RESAFE-RNN demonstrates significant improvements equivalent to a 93% reduction in the number of continuous-time constraint violations using the approximate NMPC. In terms of the magnitude of those violations, a reduction of 99.94% is calculated.
@@ IV-D Results and explainability for autonomous racing
The SHAP analysis reveals key insights into the OL solution particularly for the first Legendre-Spline coefficients as illustrated in Figure 4, for the steering angle command $\delta$ and normalized acceleration or throttle command $t_{r}\in[-1,1]$ that combines the throttle and brake commands into one variable. As the NMPC solves for the steering rate $\dot{\delta}$ and normalized throttle rate $\dot{t}_{r}$ as control actions $u(t)$, the current steering angle $input\_steer=\delta$ is fed as a parameter to the NMPC and approximate NMPC in the initial state estimation. For the steering trajectory, $\alpha^{\delta}_{1,0}$ shows the strongest explainability with the current steering measurement as in Figures 4 (a). This is expected as the first coefficient of the sequence holds the zero-order information about the time-series which are mainly influenced by the bias term in the spline. Notable patterns include increased steering under braking conditions (low feature value of $input\_throttle$, tending to -1). Moreover, the velocity $input\_vx$ has major impact on the output of the approximate NMPC for $\alpha^{\delta}_{1,0}$. The high values of steering $\alpha^{\delta}_{1,0}$ (both positive and negative) occur at lower velocities. This aligns with the expected behavior, as the NMPC minimizes steering at high-speeds of over 300kph to maintain path stability. Due to orthogonality, the steering rate (Figures 4 (b)) is mainly represented by the second coefficient of the sequence, $\alpha^{\delta}_{1,1}$ which carries information on the first-order derivative with respect to time. It correlates strongly with the yaw rate, showing a compensation behavior: negative yaw rates (blue or low feature value of $input\_yawrate$) trigger positive steering rates and vice versa. That is, when the vehicle is rotating counterclockwise, the NMPC reacts by steering clockwise and vice versa. Path deviations ($input\_w$) also influence steering rates, with leftward deviations (red or high feature value) triggering a clockwise (negative) corrections and steering rates. A deceleration maneuver (negative $input\_throttle$ in blue) causes higher steering rates, mainly as the vehicle attacks corners at reduced speeds. The predominance of left-hand corners in the racing track is reflected in the asymmetric distribution of the SHAP values in steering and steering rates towards the positive right-hand side of the plot.
For throttle control, $\alpha^{t_{r}}_{1,0}$ is impacted by the current normalized throttle measurement $input\_throttle$. In fact, the OCP solves for the TLS at the initial time $t=t_{0}$ to be equal to $input\_throttle$ and this is clearly illustrated in Figure 4 (c). Moreover, $\alpha^{t_{r}}_{1,1}$ reveals the NMPC’s lap-time optimization strategy. High throttle rates (right hand of the SHAP plot in Figures 4 (d)) occur primarily at low speeds or following braking maneuvers (low feature value of $input\_vx$ and $input\_throttle$ respectively). This indicates the controller’s tendency to maximize the acceleration for minimal lap time while maintaining the vehicle stability. The orthogonality of Legendre-Spline coefficients enables this clear separation between zero-order behavior and dynamic responses, providing interpretable insights into the NMPC’s decision-making process.
@@ V AI as performance monitor for the NMPC
After demonstrating the approximate NMPC, we focus on the use of explainable performance monitors for system level KPIs monitoring as will be presented in this section. The importance of employing XAI tools such as SHAP and SR are also highlighted as we extract important insights on the CL operational capability of the NMPC by relying on a small dataset.
@@ V-A Explainable AI for performance prediction
An Explainable NMPC aids users in visually understanding key parameters influencing the decision-making process, enabling performance real-time monitoring and suggesting when fallback controllers are necessary. As shown in Figure 1, several instantaneous CL KPIs can be monitored and forecasted before occurring. We focus on two KPIs: the NMPC optimal cost-function $K_{1}(t)$ indicating optimization feasibility and system energy, and the NMPC execution time $K_{2}(t)$ reflecting real-time computation capabilities.
@@ V-B Symbolic performance monitor
We train MORRF and SR for performance prediction on $K_{1},K_{2}$. In general, the performance prediction using MORRF achieves superior accuracy and faster training (MSE: 1.8e-04, quasi-instantaneous) compared to SR (MSE: 3.3e-03, couple of minutes). This indicates that MORRF is able to capture better the complex coupling between those KPIs and the input features. However, SR provides explicit models linking $K_{1},K_{2}$ to the input features by optimizing for both the structure and parameters of the model as shown in the explainability block of Figure 1. This can enable cluster creation, and output reverse engineering. That is, if a desired computation time is to be met, an analytical operational domain of the input features can be computed by using the equations from PySR (Figure 1). Furthermore, one could employ this approach to reverse engineer the designed cost function of an operating blackbox NMPC by relying on the initial state conditions to imitate the NMPC’s internal optimization.
@@ V-C Results and analysis
Analysis of three NMPC tuning in the AVP demonstrator reveal interesting patterns for reverse engineering the NMPC cost function. In Figure 5, the first and baseline tuning indicates high sensitivity to velocity tracking error $input\_refv\_error$, while path deviations $input\_w$ to the right of the path (negative values as low feature values in blue) impact the cost more than left-hand ones (red). Increasing the velocity tracking error $e_{v}$ weight from 3.1 to 20 amplifies the velocity error component’s influence on the monitor prediction in the second tuning, as shown by the SHAP plot. Moreover, braking maneuvers (low feature values in blue of $input\_throttle$) also cause high cost, as they naturally lead to a slower velocity and a higher tracking error. Increasing the lateral tracking error weights on path deviation $w$ and heading deviation $\theta$ in the third tuning, makes the path deviation $input\_w$ the dominant feature in the explained monitor, exhibiting the expected quadratic cost behavior. In fact, highly positive and highly negative feature values of $input\_w$ (red and blue) increase the predicted NMPC cost and middle values (around zero) decrease it.
@@
Figure 5: Performance prediction on the cost function KPI: reverse engineering different controller tuning on an AVP use case
@@
Explainability of the computation time prediction for the AVP demonstrator reveals four dominant features as in Figure 6: velocity tracking error ($input\_refv\_error$), path deviation ($input\_w$), obstacle position and heading in the frame tangent to the path ($input\_paramCBF\_*obs$), and the normalized throttle command ($input\_throttle$). Large velocity tracking errors (red) significantly increase computation time, while deceleration and braking commands (low feature value, blue) demand more computational resources than acceleration and throttling commands (high feature value, red), suggesting potential numerical challenges of the NMPC to solve at low speeds as the NMPC is more sensitive to braking ($t_{r}=-1$) than accelerating ($t_{r}=+1$).
@@
Figure 6: Performance monitor and prediction: computation time in an AVP
@@
Furthermore, we run the performance monitor for the autonomous racing demonstrator and use SHAP to understand the edge cases, as in Figure 7. MORRF effectively captures both the NMPC’s optimal CL cost function value and computation time KPIs. While the monitor accurately predicts significant cost function fluctuation as for e.g. around $t=$ 90 seconds, the absolute computation time values are hardware-dependent and might be less generalizable. Yet, MORRF’s capability to handle outliers, allows it to detect sudden computational peaks proving its crucial importance for the system safety monitoring. Those peaks, although rather limited, are important for edge cases studies and understanding NMPC’s handling near the limits. A critical instance occurs near $t=$ 45 seconds, where the vehicle exits the apex of a tight corner at 60kph before transitioning to an acceleration phase towards 330kph, as predicted by the NMPC over the next 7 seconds. SHAP analysis of this instance reveals that yaw rate $input\_yawrate$ and path heading deviation $input\_theta$ are the primary contributors to the well predicted high computation time. This insight provides control engineers with three actionable options: controller redesign to better handle high yaw rate scenarios, investigation of the bicycle model’s numerical behavior during the optimization under large yaw rate, or implementation of a fallback controller during such challenging cornering scenarios. Further analysis at the high-speed chicane (near $t=$ 125 seconds) reveal more insights on the computation time patterns. While some instances maintain normal execution times around the mean, critical cases emerge from the combined effects of large normalized throttle (i.e. acceleration) and yaw rate. As shown in the DT snapshot of Figure 7, the NMPC commands a 50% full throttle, and this occurs simultaneously under large rotation or yaw rates, as depicted in the SHAP plot. This causes a sudden surge in computation time possibly due to the yaw rate and non-slip constraints being activated. This dynamic corner scenario, with its rapidly varying NMPC solutions, likely impacts the OSQP solver’s active-set method, particularly when reusing the previous iteration’s solutions as warm-start. Finally, the complete execution time monitor’s explainability plot is presented in Figure 1.

### v2

@@ III-B MultiOutput Regression Trees, Neural Networks and Symbolic Regression
Finally, SR is another regression model that is explainable by design. SR offers interpretable analytical equations linking outputs to inputs, providing physical insight to the control engineer through explicit mathematical formulas. Moreover, the approach learns a structure of the underlying physics if the used basis functions capture the pattern well. We use PySR [1] with the binary operators {++, -, $\times$, $\operatorname{atan2}(y,x)$} and unary operators $\cos(x)$, $\sin(x)$, $\exp(x)$, $|x|$, $x^{2}$}, with maximum 200 iterations and 500 cycles per iteration. We apply SR specifically for KPIs prediction and monitoring, where the output dimensionality is manageable.
III-C Data generation: autonomous driving and racing
We demonstrate our work on autonomous driving and racing control applications. The NMPC is implemented in CasADi [19] using an SQP method with OSQP as the underlying QP solver. The OCP is transcribed into a nonlinear programming problem using RESAFE/COL [16]. The NMPC uses a fused kinematic-dynamics bicycle model with a Pacejka tire formulation. The verification uses 15 Degrees-of-Freedom high-fidelity Digital Twin (DT) of the vehicles in Simcenter Amesim. Two demonstration setups are created:
@@ III-C Data generation: autonomous driving and racing
Both setups integrate the NMPC as a standalone C-code library for co-simulation with the Digital Twins (DTs) with Simcenter Amesim for vehicle dynamics and Simcenter Prescan for environment simulation and sensor modeling for obstacle detection of crossing pedestrians and road users as visualized in Figure 7. The NMPC serves as the lowest-level control, executing trajectories from a high-level planner through steering, brake and throttle commands. The collected data is divided into training (64%), validation (16%) and testing (20%). Finally, the data is normalized per feature and per output to [1,1][-1,1].
@@ IV AI as an explainable NMPC approximation
IV-A RESAFE/COL for physics-informed approximate NMPC
The NMPC solves for continuous-time trajectories that are embedded in form of Legendre-Splines’ coefficients (3). As explained in Sec. II-A, the use of a Legendre-Spline with orthogonal basis offers two key advantages: 1) naturally regularized low-order embedding through decoupled and independent coefficients, and 2) physical constraint enforcement via a linear coefficient mapping, enabling CTCP without discrete sampling of the time-series trajectory.
As illustrated in Figure 1, we collect CL data of the NMPC controlling a high-fidelity DT to train ML regression models to imitate the NMPC’s OL solution. Using similar input parameters being fed into the NMPC, the trained regression model would mimic the optimization process. Unlike standard imitation learning approaches that learn only the NMPC’s policy or first control action $\bm{\mathrm{u}}(t=t_{0})$, we propose to learn the complete time-series of this policy evolution $\bm{\mathrm{u}}_{(\cdot)}(t)$, and that of the states $\bm{\mathrm{x}}_{(\cdot)}(t)$. This approach captures the optimization framework linking the predicted states and control actions. It also enables an effective warm-starting strategy for the NMPC which is known to have benefits on the numerical efficiency and helps speeding-up the computation in methods such as SQP. A baseline approach to this OL trajectory regression predicts the sequence of the sampled discrete trajectory points, or the embedding coefficients using a Mean Squared Error (MSE) loss:
$\mathcal{L}=\mathcal{L}_{MSE}=\sum_{i=1}^{N_{batch}}\lVert(\bar{\bm{\alpha}}_{i}-\bm{\alpha}_{i})\rVert^{2}/N_{batch},$ (5)
@@ IV-A RESAFE/COL for physics-informed approximate NMPC
where $\epsilon_{tol}$ defines the violation threshold tolerance. Note that linear state and control constraints of the form $\underline{\bm{\mathrm{x}}}\leq\bm{\mathrm{x}}(t)\leq\overline{\bm{\mathrm{x}}}$ are expressed in the generic form of $g(\bm{\mathrm{x}},\bm{\mathrm{u}})\leq 0$ for conciseness. The extrema elements of the convex hulls $\mathcal{P}_{u}$ and $\mathcal{P}_{v_{x}}$ are shown in the shaded yellow areas in Figures 1 and 3, over the decoded control actions and velocity trajectories.
@@ IV-B Multistep prediction using coefficients: a data efficient and explainable approach
The proposed Legendre-Spline embedding addresses the challenges stated in Sec. II-A by encoding physical information through the coefficients: the zero order coefficient $\alpha_{0}$ for the trajectory bias or mean term, the first order coefficient $\alpha_{1}$ for information about the rate of change, the second-order coefficient $\alpha_{2}$ about acceleration characteristics and the higher-order terms on additional dynamic features of the trajectory. This representation enables engineers to interpret and shape the OCP’s OL behavior through physically meaningful parameters that can be explained using SHAP. Therefore, by combining (3) with (6) and XAI tools, we allow multistep sequence prediction in one shot while requiring little data due to the low dimensionality of the prediction, and while having physical insights as shown in the framework of Figure 1.
@@ IV-C ExAMPC as a warm-starter for NMPC
We train two RNNs to approximate the NMPC for the AVP use case: one using the baseline with MSE (MSE-RNN) and another with the RESAFE/COL-type of loss as in (6) for CTCP with $\gamma=1$ (RESAFE-RNN). As shown in Figure 2, the approximate NMPC as RESAFE-RNN achieves 556 continuous-time constraint violations out of 57632 testing instances. The total loss is equal to 2.05e-04 divided into a MSE of 2.0e-04 and CTCP of 4.9e-06. In contrast, the baseline MSE-RNN using only coefficients learning results in 8113 violations out of 57632 instances, with a MSE on the coefficients of 1.7e-04 but a CTCP of 8.5e-03. Overall, the RESAFE/COL approach with RESAFE-RNN demonstrates significant improvements equivalent to a 93% reduction in the number of continuous-time constraint violations using the approximate NMPC. In terms of the magnitude of those violations, a reduction of 99.94% is calculated. As the approximate NMPC with RESAFE-RNN 1) accurately predicts the solution of the NMPC, and 2) exhibits little constraints violations, it could be used as an initial feasible guess for the online NMPC.
@@ IV-D Results and explainability for autonomous racing
The SHAP analysis reveals key insights into the OL solution particularly for the first Legendre-Spline coefficients as illustrated in Figure 4, for the steering angle command $\delta$ and normalized acceleration or throttle command $t_{r}\in[-1,1]$ that combines the throttle and brake commands into one variable. SHAP values on the x-axis show each feature’s impact on the model’s prediction relative to the baseline (average). Positive values (right) push predictions higher, and negative values (left) lower them, with a magnitude equal to the SHAP value. Influential and important features are ordered top to bottom. Each dot represents a data point, colored, by the feature value (red for high, blue for low). As the NMPC solves for the steering rate $\dot{\delta}$ and normalized throttle rate $\dot{t}_{r}$ as control actions $\bm{\mathrm{u}}(t)$, the current steering angle $input\_steer=\delta$ is fed as a parameter to the NMPC and approximate NMPC in the initial state estimation. For the steering trajectory, $\alpha^{\delta}_{1,0}$ shows the strongest explainability with the current steering measurement as in Figures 4 (a). This is expected as the first coefficient of the sequence holds the zero-order information about the time-series which are mainly influenced by the bias term in the spline, and the OCP (1) solves for the Legendre-Spline of the steering trajectory to be equal to input_steerinput\_steer at $t=0$. Notable patterns include increased steering under braking conditions (low feature value of input_throttleinput\_throttle, tending to -1). Moreover, the velocity input_vxinput\_vx has major impact on the output of the approximate NMPC for $\alpha^{\delta}_{1,0}$. The high values of steering $\alpha^{\delta}_{1,0}$ (both positive and negative) occur at lower velocities. This aligns with the expected behavior, as the NMPC minimizes steering at high-speeds of over 300kph to maintain path stability. Due to orthogonality, the steering rate (Figures 4 (b)) is mainly represented by the second coefficient of the sequence, $\alpha^{\delta}_{1,1}$ which carries information on the first-order derivative with respect to time. It correlates strongly with the yaw rate, showing a compensation behavior: negative yaw rates (blue or low feature value of input_yawrateinput\_yawrate) trigger positive steering rates and vice versa. That is, when the vehicle is rotating counterclockwise, the NMPC reacts by steering clockwise and vice versa. Path deviations (input_winput\_w) also influence steering rates, with leftward deviations (red or high feature value) triggering a clockwise (negative) corrections and steering rates. A deceleration maneuver (negative input_throttleinput\_throttle in blue) causes higher steering rates, mainly as the vehicle attacks corners at reduced speeds. The predominance of left-hand corners in the racing track is reflected in the asymmetric distribution of the SHAP values in steering and steering rates towards the positive right-hand side of the plot. The orthogonality of Legendre-Spline coefficients enables this clear separation between zero-order behavior and dynamic responses, providing interpretable insights into the NMPC’s decision-making process.
@@ V AI as performance monitor for the NMPC
After demonstrating the approximate NMPC, we focus in this section on the use of explainable performance monitors for system level KPIs monitoring. The importance of employing XAI tools such as SHAP and SR are also highlighted as we extract important insights on the NMPC’s CL operational capability by relying on a small dataset.
@@ V-A Explainable AI for performance prediction
An Explainable NMPC aids users in visually understanding key parameters influencing the decision-making process, enabling real-time monitoring of the performance and suggesting when fallback controllers are necessary. As shown in Figure 1, several instantaneous CL KPIs can be monitored and forecast before occurring. We focus on two KPIs: the NMPC optimal cost-function $K_{1}(t)$ indicating optimization feasibility and system energy, and the NMPC execution time $K_{2}(t)$ reflecting real-time computation capabilities.
@@ V-B Symbolic performance monitor
We train MORRF and SR for performance prediction on $K_{1},K_{2}$. In general, the performance prediction using MORRF achieves superior accuracy and faster training (MSE: 1.8e-04, quasi-instantaneous) compared to SR (MSE: 2.3e-03, couple of minutes). This indicates that MORRF is able to capture better the complex coupling between those KPIs and the input features. However, SR provides explicit models linking $K_{1},K_{2}$ to the input features by optimizing for both the structure and parameters of the model as shown in the explainability block of Figure 1. This can enable cluster creation, and output reverse engineering. That is, if a desired computation time is to be met, an analytical operational domain of the input features can be computed by using the equations from PySR (Figure 1). Furthermore, one could employ this approach to reverse engineer the designed cost function of an operating blackbox NMPC by relying on the measured features to imitate the NMPC’s internal optimization.
@@ V-C Results and analysis
Analysis of three NMPC tuning in the AVP demonstrator reveal interesting patterns for reverse engineering the NMPC cost function. In Figure 5, the first (baseline) tuning indicates high sensitivity to velocity tracking error input_refv_errorinput\_refv\_error, and asymmetric responses to the path deviations input_winput\_w with deviations to the right of the path (low-feature values) impacting the cost function more than the left-hand ones (high-feature values). Increasing the velocity tracking error weight $e_{v}$ from 3.1 to 20 amplifies the velocity error component’s influence on the monitor prediction in the second tuning, confirming the method’s ability to capture the internal NMPC optimization priorities. Increasing the lateral tracking error weights on path deviation $w$ and heading deviation $\theta$ in the third tuning, produces the expected quadratic cost behavior with input_winput\_w becoming a dominant feature in the explainable monitor, where both high and low-feature values increase the predicted NMPC cost or the consumed energy.
@@
Figure 5: Performance prediction on the cost function KPI: reverse engineering different controller tuning in an AVP use case
@@
Explainability of the computation time prediction for the AVP demonstrator reveals four dominant features as in Figure 6: velocity tracking error (input_refv_errorinput\_refv\_error), path deviation (input_winput\_w), obstacle position and heading in the frame tangent to the path (input_paramCBF_obsinput\_paramCBF\_*obs), and the normalized throttle command (input_throttleinput\_throttle). Large velocity tracking errors (red) significantly increase computation time, while deceleration and braking commands (low feature value, blue) demand more computational resources than acceleration and throttling commands (high feature value, red), suggesting potential numerical challenges of the NMPC to solve at low speeds as the NMPC is more sensitive to braking ($t_{r}=-1$) than accelerating ($t_{r}=+1$).
@@
Figure 6: Explainable performance monitor and prediction for the NMPC computation time in an AVP use case: isolating the dominant features
@@
Furthermore, we run the performance monitor for the autonomous racing demonstrator and use SHAP to understand the edge cases, as in Figure 7. MORRF effectively captures both the NMPC’s optimal CL cost function value and computation time KPIs. While the monitor accurately predicts significant cost function fluctuation as for e.g. around $t=$ 90s, the absolute computation time values are hardware-dependent and might be less generalizable. Yet, MORRF’s capability to handle outliers, allows it to detect sudden computational peaks proving its crucial importance for the system safety monitoring. Those peaks, although rather limited, are important for edge cases studies and understanding NMPC’s handling near the limits. A critical instance occurs near $t=$ 45s, where the vehicle exits the apex of a tight corner at 60kph before transitioning to an acceleration phase towards 330kph, as predicted by the NMPC over the next 7s. SHAP analysis of this instance reveals that yaw rate input_yawrateinput\_yawrate and path heading deviation input_thetainput\_theta are the primary contributors to the well predicted high computation time. This insight provides control engineers with three actionable options: controller redesign to better handle high yaw rate scenarios, investigation of the bicycle model’s numerical behavior during the optimization under large yaw rate, or implementation of a fallback controller during such challenging cornering scenarios. Further analysis at the high-speed chicane (near $t=$ 125s) reveal more insights on the computation time patterns. While some instances maintain normal execution times around the mean, critical cases emerge from the combined effects of large normalized throttle (i.e. acceleration) and yaw rate. As shown in the DT snapshot of Figure 7, the NMPC commands a 50% full throttle, and this occurs simultaneously under large rotation or yaw rates, as depicted in the SHAP plot. This causes a sudden surge in computation time. The sudden activation of yaw rate and non-slip constraints in this dynamic corner scenario alters the optimal solution. This makes the warm-started primal and dual variables from the previous iteration a less effective initial guess for the new problem instance, thereby increasing the number of OSQP iterations. Finally, the complete execution time monitor’s explainability plot is presented in Figure 1.